\begin{thebibliography}{DGMS75}
\newcommand{\enquote}[1]{``#1''}
\providecommand{\url}[1]{\texttt{#1}}
\providecommand{\urlprefix}{URL }
\providecommand{\bibinfo}[2]{#2}
\providecommand{\eprint}[2][]{\url{#2}}

\bibitem[ABR92]{ABR}
\bibinfo{author}{D.~Arapura}, \bibinfo{author}{P.~Bressler} \& \bibinfo{author}{M.~Ramachandran}.
\newblock \enquote{\bibinfo{title}{On the fundamental group of a compact {K}\"ahler manifold}}.
\newblock \emph{\bibinfo{journal}{Duke Math. J.}}, \textbf{\bibinfo{volume}{68}(\bibinfo{year}{1992})(\bibinfo{number}{3})}:\bibinfo{pages}{477--488}.
\newblock \urlprefix\url{http://dx.doi.org/10.1215/S0012-7094-92-06819-0}.

\bibitem[Bau62]{Baumslag}
\bibinfo{author}{G.~Baumslag}.
\newblock \enquote{\bibinfo{title}{On generalised free products}}.
\newblock \emph{\bibinfo{journal}{Math. Z.}}, \textbf{\bibinfo{volume}{78}(\bibinfo{year}{1962})}:\bibinfo{pages}{423--438}.
\newblock \urlprefix\url{http://dx.doi.org/10.1007/BF01195185}.

\bibitem[BNS87]{BNS}
\bibinfo{author}{R.~Bieri}, \bibinfo{author}{W.~D. Neumann} \& \bibinfo{author}{R.~Strebel}.
\newblock \enquote{\bibinfo{title}{A geometric invariant of discrete groups}}.
\newblock \emph{\bibinfo{journal}{Invent. Math.}}, \textbf{\bibinfo{volume}{90}(\bibinfo{year}{1987})(\bibinfo{number}{3})}:\bibinfo{pages}{451--477}.
\newblock \urlprefix\url{http://dx.doi.org/10.1007/BF01389175}.

\bibitem[BS76]{BuanicuaStuanuac}
\bibinfo{author}{C.~B{\u{a}}nic{\u{a}}} \& \bibinfo{author}{O.~St{\u{a}}n{\u{a}}{\c{s}}il{\u{a}}}.
\newblock \emph{\bibinfo{title}{Algebraic methods in the global theory of complex spaces}}.
\newblock \bibinfo{publisher}{Editura Academiei, Bucharest; John Wiley \& Sons, London-New York-Sydney} (\bibinfo{year}{1976}).
\newblock \bibinfo{note}{Translated from the Romanian}.

\bibitem[Cam95]{Campana}
\bibinfo{author}{F.~Campana}.
\newblock \enquote{\bibinfo{title}{Remarques sur les groupes de {K}\"ahler nilpotents}}.
\newblock \emph{\bibinfo{journal}{Ann. Sci. \'Ecole Norm. Sup. (4)}}, \textbf{\bibinfo{volume}{28}(\bibinfo{year}{1995})(\bibinfo{number}{3})}:\bibinfo{pages}{307--316}.
\newblock \urlprefix\url{http://www.numdam.org/item?id=ASENS_1995_4_28_3_307_0}.

\bibitem[{Cla}08]{Claudon}
\bibinfo{author}{B.~{Claudon}}.
\newblock \enquote{\bibinfo{title}{{Convexit{\'e} holomorphe du rev{\^e}tement de Malcev d'apr{\`e}s S. Leroy}}}.
\newblock \emph{\bibinfo{journal}{arXiv e-prints}}, \textbf{(\bibinfo{year}{2008})}:\bibinfo{eid}{arXiv:0809.0920}.
\newblock \eprint{0809.0920}, \urlprefix\url{http://dx.doi.org/10.48550/arXiv.0809.0920}.

\bibitem[Del10]{Delzant10}
\bibinfo{author}{T.~Delzant}.
\newblock \enquote{\bibinfo{title}{L'invariant de {B}ieri-{N}eumann-{S}trebel des groupes fondamentaux des vari\'et\'es k\"ahl\'eriennes}}.
\newblock \emph{\bibinfo{journal}{Math. Ann.}}, \textbf{\bibinfo{volume}{348}(\bibinfo{year}{2010})(\bibinfo{number}{1})}:\bibinfo{pages}{119--125}.
\newblock \urlprefix\url{http://dx.doi.org/10.1007/s00208-009-0468-8}.

\bibitem[DGMS75]{DGMS}
\bibinfo{author}{P.~Deligne}, \bibinfo{author}{P.~Griffiths}, \bibinfo{author}{J.~Morgan} \& \bibinfo{author}{D.~Sullivan}.
\newblock \enquote{\bibinfo{title}{Real homotopy theory of {K}\"ahler manifolds}}.
\newblock \emph{\bibinfo{journal}{Invent. Math.}}, \textbf{\bibinfo{volume}{29}(\bibinfo{year}{1975})(\bibinfo{number}{3})}:\bibinfo{pages}{245--274}.
\newblock \urlprefix\url{http://dx.doi.org/10.1007/BF01389853}.

\bibitem[Gro89]{Gromov}
\bibinfo{author}{M.~Gromov}.
\newblock \enquote{\bibinfo{title}{Sur le groupe fondamental d'une vari\'et\'e{} k\"ahl\'erienne}}.
\newblock \emph{\bibinfo{journal}{C. R. Acad. Sci. Paris S\'er. I Math.}}, \textbf{\bibinfo{volume}{308}(\bibinfo{year}{1989})(\bibinfo{number}{3})}:\bibinfo{pages}{67--70}.

\bibitem[Hai87]{Hain87}
\bibinfo{author}{R.~M. Hain}.
\newblock \enquote{\bibinfo{title}{Higher {A}lbanese manifolds}}.
\newblock \enquote{\bibinfo{booktitle}{Hodge theory ({S}ant {C}ugat, 1985)}}, vol. \bibinfo{volume}{1246} of \emph{\bibinfo{series}{Lecture Notes in Math.}}, \bibinfo{pages}{84--91}. \bibinfo{publisher}{Springer, Berlin} (\bibinfo{year}{1987}):\hspace{0pt}\urlprefix\url{http://dx.doi.org/10.1007/BFb0077531}.

\bibitem[Kat97]{Katzarkov97}
\bibinfo{author}{L.~Katzarkov}.
\newblock \enquote{\bibinfo{title}{Nilpotent groups and universal coverings of smooth projective varieties}}.
\newblock \emph{\bibinfo{journal}{J. Differential Geom.}}, \textbf{\bibinfo{volume}{45}(\bibinfo{year}{1997})(\bibinfo{number}{2})}:\bibinfo{pages}{336--348}.
\newblock \urlprefix\url{http://projecteuclid.org/euclid.jdg/1214459801}.

\bibitem[KR98]{KatzarkovRamachandran98}
\bibinfo{author}{L.~Katzarkov} \& \bibinfo{author}{M.~Ramachandran}.
\newblock \enquote{\bibinfo{title}{On the universal coverings of algebraic surfaces}}.
\newblock \emph{\bibinfo{journal}{Ann. Sci. \'{E}cole Norm. Sup. (4)}}, \textbf{\bibinfo{volume}{31}(\bibinfo{year}{1998})(\bibinfo{number}{4})}:\bibinfo{pages}{525--535}.
\newblock \urlprefix\url{http://dx.doi.org/10.1016/S0012-9593(98)80105-5}.

\bibitem[Ler98]{Leroy}
\bibinfo{author}{S.~Leroy}.
\newblock \enquote{\bibinfo{title}{Vari\'et\'es d'{A}lbanese sup\'erieures d'une vari\'et\'e{} k\"ahl\'erienne compacte}}.
\newblock \emph{\bibinfo{journal}{C. R. Acad. Sci. Paris S\'er. I Math.}}, \textbf{\bibinfo{volume}{327}(\bibinfo{year}{1998})(\bibinfo{number}{3})}:\bibinfo{pages}{295--300}.
\newblock \urlprefix\url{http://dx.doi.org/10.1016/S0764-4442(98)80149-3}.

\bibitem[Liu23]{Yuan23}
\bibinfo{author}{Y.~Liu}.
\newblock \enquote{\bibinfo{title}{On the holomorphic convexity of reductive {G}alois coverings over compact {K}\"{a}hler surfaces}}.
\newblock \emph{\bibinfo{journal}{Adv. Math.}}, \textbf{\bibinfo{volume}{431}(\bibinfo{year}{2023})}:\bibinfo{pages}{Paper No. 109245, 19}.
\newblock \urlprefix\url{http://dx.doi.org/10.1016/j.aim.2023.109245}.

\bibitem[LL99]{LeroyThesis}
\bibinfo{author}{S.~Leroy-Leli{\`e}vre}.
\newblock \emph{\bibinfo{title}{{Vari{\'e}t{\'e}s d'Albanese sup{\'e}rieures complexes d'une vari{\'e}t{\'e} kahl{\'e}rienne compacte}}}.
\newblock \bibinfo{type}{Theses}, \bibinfo{school}{{Universit{\'e} Henri Poincar{\'e} - Nancy 1}} (\bibinfo{year}{1999}).
\newblock \bibinfo{note}{Texte int{\'e}gral accessible uniquement aux membres de l'Universit{\'e} de Lorraine}, \urlprefix\url{https://hal.univ-lorraine.fr/tel-01747060}.

\bibitem[Mal49]{Malcev}
\bibinfo{author}{A.~I. Mal'tsev}.
\newblock \enquote{\bibinfo{title}{On a class of homogeneous spaces.}}
\newblock \emph{\bibinfo{journal}{Izv. Akad. Nauk SSSR, Ser. Mat.}}, \textbf{\bibinfo{volume}{13}(\bibinfo{year}{1949})}:\bibinfo{pages}{9--32}.

\bibitem[NR95]{NR95}
\bibinfo{author}{T.~Napier} \& \bibinfo{author}{M.~Ramachandran}.
\newblock \enquote{\bibinfo{title}{Structure theorems for complete {K}\"ahler manifolds and applications to {L}efschetz type theorems}}.
\newblock \emph{\bibinfo{journal}{Geom. Funct. Anal.}}, \textbf{\bibinfo{volume}{5}(\bibinfo{year}{1995})(\bibinfo{number}{5})}:\bibinfo{pages}{809--851}.
\newblock \urlprefix\url{http://dx.doi.org/10.1007/BF01897052}.

\bibitem[NR01]{NR01GAFA}
---{}---{}---.
\newblock \enquote{\bibinfo{title}{Hyperbolic {K}\"ahler manifolds and proper holomorphic mappings to {R}iemann surfaces}}.
\newblock \emph{\bibinfo{journal}{Geom. Funct. Anal.}}, \textbf{\bibinfo{volume}{11}(\bibinfo{year}{2001})(\bibinfo{number}{2})}:\bibinfo{pages}{382--406}.
\newblock \urlprefix\url{http://dx.doi.org/10.1007/PL00001677}.

\bibitem[Uen75]{Ueno}
\bibinfo{author}{K.~Ueno}.
\newblock \emph{\bibinfo{title}{Classification theory of algebraic varieties and compact complex spaces}}, vol. \bibinfo{volume}{Vol. 439} of \emph{\bibinfo{series}{Lecture Notes in Mathematics}}.
\newblock \bibinfo{publisher}{Springer-Verlag, Berlin-New York} (\bibinfo{year}{1975}).
\newblock \bibinfo{note}{Notes written in collaboration with P. Cherenack}.

\end{thebibliography}

\documentclass[11pt,english]{smfart} 
\usepackage[T1]{fontenc}
\usepackage[english,french]{babel}
\usepackage{amscd}
\usepackage{amssymb,url,xspace,smfthm}
\usepackage{amsbsy,bm} 
\usepackage{paralist}
\usepackage{datetime} 
\usepackage{colortbl}
\usepackage{color}
\usepackage[dvipsnames]{xcolor}
\usepackage{mathrsfs}  
\usepackage[colorlinks,	colorlinks,
linkcolor=BrickRed,
citecolor=Green,
urlcolor=Cerulean,hypertexnames=true]{hyperref} 
\usepackage{mathtools}
\usepackage{euscript}
\usepackage[all]{xy}
\usepackage{hyperref}
\usepackage{graphicx} 
\usepackage[text={150mm,251mm},centering, marginparwidth=75pt]{geometry} 
\usepackage{comment}  
\usepackage{cite}  
\usepackage{thmtools}
\usepackage{enumitem}
\usepackage{letltxmacro} 
\usepackage{nameref}
\usepackage{cleveref}
\usepackage{calc}
\usepackage{interval}
\usepackage{manfnt}
\usepackage{tikz-cd}  
\usepackage[utf8]{inputenc}
\usepackage{marginnote}
\usepackage{smfhyperref}
\usepackage[hyperpageref]{backref}
\renewcommand{\backref}[1]{$\uparrow$~#1}

\usepackage{faktor} 
\usepackage{scalerel}

\makeatletter     
\newcommand{\crefnames}[3]{%
	\@for\next:=#1\do{%
		\expandafter\crefname\expandafter{\next}{#2}{#3}%
	}%
}
\makeatother

\setlist[itemize]{wide = 0pt, labelwidth = 2em, labelsep*=0em, itemindent = 0pt, leftmargin = \dimexpr\labelwidth + \labelsep\relax, noitemsep,topsep = 1ex,}
\setlist[enumerate]{wide = 0pt, labelwidth = 2em, labelsep*=0em, itemindent = 0pt, leftmargin = \dimexpr\labelwidth + \labelsep\relax, noitemsep,topsep = 1ex}

\theoremstyle{plain}
\newtheorem{thmx}{Theorem}
\renewcommand{\thethmx}{\Alph{thmx}} 
\newtheorem{theorem}{Theorem}[section]  
\newtheorem{lemma}[theorem]{Lemma}
 
\newtheorem{proposition}[theorem]{Proposition}

\theoremstyle{definition}
 
\theoremstyle{remark}
\newtheorem{remark}[theorem]{Remark}

\numberwithin{equation}{section}  

\theoremstyle{plain}
\newlist{thmlist}{enumerate}{1}
\setlist[thmlist]{wide = 0pt, labelwidth = 2em, labelsep*=0em, itemindent = 0pt, leftmargin = \dimexpr\labelwidth + \labelsep\relax, noitemsep,topsep = 1ex, font=\normalfont, label=(\roman*), ref=\thetheorem.(\roman{thmlisti})}

\addtotheorempostheadhook[theorem]{\crefalias{thmlisti}{thm}}

\addtotheorempostheadhook[assumpsion]{\crefalias{thmlisti}{assumption}}

\addtotheorempostheadhook[corollary]{\crefalias{thmlisti}{cor}}

\addtotheorempostheadhook[proposition]{\crefalias{thmlisti}{proposition}}

\addtotheorempostheadhook[definition]{\crefalias{thmlisti}{dfn}}

\addtotheorempostheadhook[lemma]{\crefalias{thmlisti}{lem}}
\addtotheorempostheadhook[main]{\crefalias{thmlisti}{main}}

\addtotheorempostheadhook[remark]{\crefalias{thmlisti}{rem}}

\newlist{thmenum}{enumerate}{1} 
\setlist[thmenum]{wide = 0pt, labelwidth = 2em, labelsep*=0em, itemindent = 0pt, leftmargin = \dimexpr\labelwidth + \labelsep\relax, noitemsep,topsep = 1ex, font=\normalfont, label=(\roman*), ref=\thethmx.(\roman{thmenumi})}
\crefalias{thmenumi}{thmx} 

\newlist{corlist}{enumerate}{1} 
\setlist[corlist]{wide = 0pt, labelwidth = 2em, labelsep*=0em, itemindent = 0pt, leftmargin = \dimexpr\labelwidth + \labelsep\relax, noitemsep,topsep = 1ex, font=\normalfont, label=(\roman*), ref=\thecorx.(\roman{corlisti})}
\crefalias{corlisti}{corx} 

\crefname{lemma}{Lemma}{Lemmas} 
\crefname{conjecture}{Conjecture}{Conjectures}
\crefname{theorem}{Theorem}{Theorems}
\crefname{proposition}{Proposition}{Propositions}
\crefname{definition}{Definition}{Definitions}
\crefname{remark}{Remark}{Remarks}
\crefname{corollary}{Corollary}{Corollaries}
\crefname{corx}{Corollary}{Corollaries}
\crefname{problem}{Problem}{Problems}
\crefname{thmx}{Theorem}{Theorems}
\crefname{claim}{Claim}{Claims}
\crefname{assumption}{Assumption}{Assumptions}
\crefname{main}{Main Theorem}{Main Theorems}

\newcommand{\C}{\mathbb{C}}
\newcommand{\R}{\mathbb{R}}

\newcommand{\Z}{\mathbb{Z}}
\newcommand{\norm}[1]{\left\lVert#1\right\rVert}
\newcommand{\pbar}{\overline{\partial}}

\begin{document}

\title[The holomorphic convexity of nilpotent coverings]{On the holomorphic convexity of nilpotent coverings over compact K\"ahler surfaces}

{
\author{Yuan Liu}
\email{yuanliu@ustc.edu.cn}
\address{Institute of Geometry and Physics, University of Science and Technology of China, Hefei, China.} 
\urladdr{https://sites.google.com/view/yuan-lius-website} 
}

\begin{abstract}
We prove that any nilpotent regular covering over a compact K\"ahler surface is holomorphically convex if it does not have two ends. Furthermore, we show that the Malcev covering of any compact K\"ahler manifold has at most one end.
\end{abstract}
\maketitle

\section{Introduction}
The Shafarevich conjecture asks if the universal covering of a compact K\"ahler manifold is holomorphically convex. Not only focusing on the universal coverings, we can also ask the same question for intermediate coverings, and find natural conditions under which these intermediate coverings are holomorphically convex. This kind of question was answered partially for the Malcev covering over projective manifolds by Katzarkov (c.f. \cite{Katzarkov97}), for reductive coverings without two ends over projective surfaces by Katzarkov and Ramachandran (c.f. \cite{KatzarkovRamachandran98}), for the Malcev covering over compact K\"ahler manifolds by Leroy (c.f. \cite{LeroyThesis}, explained by Claudon in \cite{Claudon}), and for reductive coverings without two ends over compact K\"ahler surfaces by the author (c.f. \cite{Yuan23}). This is a continuation about considering the virtually nilpotent coverings over a compact K\"ahler surface and we study when they are holomorphically convex. 

Recall that for any regular covering over a compact K\"ahler manifold, we call it a $\mathcal{P}$-covering if the Galois group of this regular covering has the property $\mathcal{P}$. The property $\mathcal{P}$ could be infinite, reductive, nilpotent, linear, etc. All the coverings considered afterward will be regular without explicitly being mentioned. Also,  we say a group is virtually $\mathcal{P}$ if it admits a normal subgroup of finite index satisfying the property $\mathcal{P}$. Notice that for holomorphic convexity, we only need to consider the infinite coverings, and a virtually $\mathcal{P}$-covering is holomorphically convex if and only if the corresponding $\mathcal{P}$-covering is.

We focus on the case of surfaces due to technical limitations (see \cref{rmk:why_two_dim}), and our first main result is:
\begin{thmx}[\cref{thm:nilpotent_covering}]
Let $\hat{X}$ be an infinite virtually nilpotent covering of a compact K\"ahler surface and it does not have two ends, then $\hat{X}$ is holomorphically convex.
\end{thmx}

Being nilpotent is the next algebraic complication after being abelian and the proof of the above theorem is inspired by that in the abelian case (\cref{prop:abelian_cover}). The condition that $\hat{X}$ does not have two ends might be weird at first glance, and we add an explanation of this in \cref{subsec:two_ends}. A criterion (\cref{thm:two_ends_BNS}) is given for the holomorphic convexity of two-ended coverings, which is not new and should be seen as a different sight of a theorem of Napier and Ramachandran using the BNS-invariant.

We also prove a result on the number of ends for the Malcev covering of a compact K\"ahler manifold.

\begin{thmx}[\cref{thm:one_end_malcev}]
The number of ends for the Malcev covering of a compact K\"ahler manifold is at most one.    
\end{thmx}

The Malcev covering can be roughly seen as the universal covering with the restriction of being nilpotent and torsion-free, and this is analog to the same result for universal coverings (\cite{Gromov},\cite{ABR}).

The main techniques we apply to prove these results are higher Albanese manifolds introduced by Hain (\cite{Hain87}, see \cref{subsec:higher_alb}) and the canonical coordinates for nilpotent groups introduced by Malcev (\cite{Malcev}, see \cref{subsec:canonical_coord}).

\section{On the holomorphic convexity of abelian coverings}

\subsection{The proof in the abelian case}
The easiest situation is for an abelian covering without two ends over a compact K\"ahler surface, which is well-known to experts (page 526, \cite{KatzarkovRamachandran98}). We include its proof here for completeness. This also inspires us for the proof in the nilpotent case.

Let $X$ be a compact K\"ahler surface with the universal covering $\widetilde{X}$, and $\Gamma=\pi_1(X)$ be its fundamental group. Take $\rho:\Gamma\to A$ as a group homomorphism onto an abelian group $A$. We take $X_{\rho}=\widetilde{X}\slash \mathrm{Ker}(\rho)$. This is the regular covering of $X$ with the properties that $\mathrm{Gal}(X_{\rho}\slash X)=A$ and $\pi_1(X_{\rho})=\mathrm{Ker}(\rho)$ and we will call $X_{\rho}$ the covering of $X$ associated with $\rho$. Since $\Gamma$ is finitely generated and $A$ is abelian, we know the torsion $\mathrm{Tor}(A)$ is a finite group. We will study the holomorphic convexity of these intermediate coverings. Since a finite ramified covering of a holomorphically convex covering is also holomorphically convex, we may assume that $A$ is torsion-free.

With the notations mentioned above, we have
\begin{proposition}\label{prop:abelian_cover}
Let ${X}_{\rho}$ be an infinite abelian covering over a compact K\"ahler surface $X$ and it does not have two ends, then ${X}_{\rho}$ is holomorphically convex.
\end{proposition}

\begin{proof}
We may assume that the Galois group is $A=\Z^r$ for some $r\geqslant 1$.
By the universal property of the Albanese map, we know that $\rho$ factors through the Albanese map of $X$, say $\alpha: X\to \mathrm{Alb}(X)$, i.e. we have the commutative diagram
\begin{center}
\begin{tikzcd}
\Gamma \ar[d,swap,"\alpha_{*}"] \ar[r,"\rho"]& A \\
\pi_1(\mathrm{Alb}(X))\ar[ru,swap,"\rho"] &
\end{tikzcd}    
\end{center}

Now take the covering of $X$(resp. $Y:=\mathrm{Alb(X)}$) associated with $\rho$, we have 

\begin{center}
\begin{tikzcd}
X_{\rho} \ar[d,swap,"p_1"] \ar[r,"\alpha_{\rho}"]& {Y}_{\rho}\ar[d,"p_2"]\\
X\ar[r,"\alpha"] & Y
\end{tikzcd}    
\end{center}
where the lifting $\alpha_{\rho}$ of $\alpha$ is holomorphic and proper.

Now we need the assumption that ${X}_{\rho}$ does not have two ends, i.e. $r>1$. Thus the regular covering $Y_{\rho}$ also does not have two ends and can be written as $Y_{\rho}=T\times \R^r$, where $T$ is some torus. Denote the composition of $\alpha_{\rho}$ with the projection onto $\R^r$ as $f=(f_1,\cdots,f_r)$. Since $f_i$'s can be seen as the real or imaginary part of the holomorphic function from $X_{\rho}$ to the universal covering of $Y$(which is a complex vector space), thus they are pluriharmonic. Let $g=\sum\limits_{i=1}^{r}f_i^2:X_{\rho}\to \R$, then $g$ is exhaustive and its Levi form
$$\sqrt{-1}\partial\pbar g=2\norm{\partial {f}_1}^2+\cdots+2\norm{\partial {f}_r}^2\geqslant 0.$$

Now we need the analysis of the degeneracy locus of this Levi form. To simplify the notation, we assume that $r=2$, and the general case is similar. Consider the complex analytic subset $E:=\{x\in {X}_{\rho}:\ (\partial {f}_1 \wedge \partial{f}_2)_x=0\}$, and we have the following two cases.

\textit{Case 1:} the subset $E$ is proper, i.e. $E\subsetneq X_{\rho}$. We have that $g$ is weakly plurisubharmonic and strongly plurisubharmonic away from a proper analytic subset $E$ (i.e. $g$ is generically strongly plurisubharmonic), and ${X}_{\rho}$ is holomorphically convex by a theorem of Narasimhan (see \cref{lem:theorem_of_narasimhan} below).

\textit{Case 2:} the Levi form of $g$ degenerates everywhere, i.e. $E={X}_{\rho}$. Now $g$ is a real analytic exhaustive function of $X_{\rho}$ with its Levi form degenerated everywhere. Denote the (singular) holomorphic foliation on $X_{\rho}$ defined by $\partial f_1$(or $\partial f_2$) as $\mathscr{F}$. For any regular value $c$ of $g$, consider its level set $\{x\in X_{\rho}:g(x)=c\}$. The degeneracy locus of Levi form of $g$ in each level set is exactly the leaves of $\mathscr{F}$. This shows that the level set of $g$ is Levi-flat. By a theorem of Napier and Ramachandran (Theorem 4.6 or Theorem 4.8, \cite{NR95}), we know that $X_{\rho}$ admits a proper holomorphic map onto a Riemann surface with connected fibers, thus is holomorphically convex.
\end{proof}

To complete the proof above, we only need to recall the following theorem of Narasimhan:
\begin{theorem}[See lemma 3.1 of \cite{KatzarkovRamachandran98} for its proof]\label{lem:theorem_of_narasimhan}
Let $M$ be a complex manifold and let $g: M \to \R$ be a continuous
plurisubharmonic exhaustion function, which is a generically strongly plurisubharmonic, i.e. being strongly plurisubharmonic away from a proper complex analytic set $A \subsetneq M$. If all noncompact irreducible
components of $A$ are Stein spaces, then $M$ is holomorphically convex.  
\end{theorem}

\begin{remark}\label{rmk:why_two_dim}
We can only deal with the two-dimensional case since we need to analyze the degeneracy locus of a Levi form, and the fact that any open Riemann surface is Stein permits us to use \cref{lem:theorem_of_narasimhan}.    
\end{remark}

\subsection{Why two ends is special?}
\label{subsec:two_ends}
Due to an example of Cousin (see Example 3.9 in \cite{NR95}), there exists a $\Z$-covering over a compact K\"ahler manifold with no non-constant holomorphic function, not to say that this covering is holomorphically convex. Thus it is crucial to exclude the case of two-ends in our arguments. This kind of phenomenon does not occur when we consider the universal coverings. This is because any group $G$ with two ends must be virtually $\Z$ and this cannot happen in the K\"ahler case due to the Hodge theory (for an analog on the Malcev covering, see \cref{thm:one_end_malcev}). But we can certainly have $\Z$ as the quotient of a K\"ahler group, thus the two-ended case appear for intermediate coverings. 

We can say some things about the holomorphic convexity of virtually $\Z$-covering over a compact K\"ahler manifold, and it is simply a re-explanation of Theroem 4.3 in \cite{NR01GAFA} using Delzant's results\cite{Delzant10} on the BNS-invariant of K\"ahler groups.

\begin{theorem}\label{thm:two_ends_BNS}
Let $X$ be a compact K\"ahler manifold with fundamental group $\Gamma$ and $\rho:\Gamma\to \Z$ be a surjective homomorphism. If $[\rho]\notin \Sigma(\Gamma)$ where $\Sigma(\Gamma)$ is the BNS-invariant of $\Gamma$\footnote{For the definition and its property, see \cite{BNS}. If $\rho:\Gamma\to \R$ is a real character, we denote its equivalent class under scalar multiplication by a positive real number as $[\rho]$. These equivalent classes are collected as $S(\Gamma)$. For a subgroup $N$, we use $S(\Gamma, N)$ to denote the subset of $S(\Gamma)$ with $\rho(N)\equiv 0$.}, then the associated covering $X_{\rho}$ admits a proper holomorphic map onto a Riemann surface with connected fibers and is thus holomorphically convex.
\end{theorem}

\begin{proof}
Due to the Theorem B1 in \cite{BNS}, we know that $N=\mathrm{Ker}(\rho)$ is finitely generated if and only if $S(\Gamma,N)=\{[\rho],[-\rho]\}\subseteq \Sigma(\Gamma)$. By the explanation of the BNS-invariant for K\"ahler groups (Theorem 1.1, \cite{Delzant10}), we know that $\Sigma(\Gamma)$ is symmetric in the sense that $\Sigma(\Gamma)=-\Sigma(\Gamma)$. Thus we have $N$ is finitely generated if and only if $[\rho]\in\Sigma(\Gamma)$. Now if $[\rho]\notin\Sigma(\Gamma)$, we have $N$ is not finitely generated, then by the proof of Theorem 4.3 in \cite{NR01GAFA} the conclusion is true.
\end{proof}

\section{From abelian case to nilpotent case}
In this section, we prove our main theorem
\begin{theorem}\label{thm:nilpotent_covering}
Let $X$ be a compact K\"ahler surface with fundamental group $\Gamma$ and $\rho:\Gamma\to D$ be a group homomorphism onto an infinite virtually nilpotent group $D$. Take
${X}_{\rho}$ be the covering associated with $\rho$ and assume that it does not have two ends, then ${X}_{\rho}$ is holomorphically convex.
\end{theorem}

Recall that a group $D$ is virtually nilpotent if it contains a (normal) subgroup of finite index which is nilpotent. Since the passing to a finite covering does not affect the holomorphic convexity, we can deal with this slightly general situation. Besides, the torsion subgroup of a finitely generated nilpotent group is finite, thus we can assume without loss of generality that $D$ is \textit{torsion-free and nilpotent}.

To prove the main theorem, we need to use the higher Albanese map by R. Hain \cite{Hain87} to replace the Albanese map. The situation is more delicate due to the group actions, and we need to analyze the intermediate covering of a higher Albanese manifold.

\subsection{Higher Albanese map and manifolds revisited}\label{subsec:higher_alb}
The readers may refer to \cite{Hain87}, \cite{Leroy}, and \cite{Claudon} for more details.

Let $\Gamma$ be any finitely generated group. Consider the descending central series of $\Gamma$:
$$\Gamma = \Gamma_1 \trianglerighteq \Gamma_2 \trianglerighteq \cdots \trianglerighteq \Gamma_n \trianglerighteq \cdots$$
where $\Gamma_{i+1}=[\Gamma,\Gamma_{i}]$ for each $i$.
We say $\Gamma$ is nilpotent if this series stops in finitely many steps, i.e. $\Gamma_{n}=\{1\}$ for some $n$. Furthermore, we say $\Gamma$ is $s$-nilpotent if $\Gamma_{s+1}=\{1\}$ and $\Gamma_{s}\neq\{1\}$. For example, a nontrivial abelian group is $1$-nilpotent. 

By construction, $\Gamma\slash \Gamma_{s+1}$ is $s$-nilpotent and its subgroup of torsion element is finite. Denote $\Gamma^{s}=(\Gamma\slash \Gamma_{s+1})\slash\mathrm{Tor}$ to be the maximal torsion-free $s$-nilponent quotient of $\Gamma$. 

Now let $X$ be a compact K\"ahler manifold with fundamental group $\Gamma$. According to Hain, we can construct the $s$-th higher albanese manifold $A_s$ and holomorphic morphisms $\alpha_s$ such that the following diagram commutes:

    \begin{equation}\label{diag:higher_alb}
    \begin{tikzcd}
        & X \ar[dd,"\alpha_s",swap]\ar[ddr,"\alpha_{s-1}"]\ar[rrrdd,"\alpha_1"]& & &\\
        &&&&\\
        \ ...\ \ar[r]&A_s \ar[r,"\pi_{s-1}"]& A_{s-1} \ar[r,"\pi_{s-2}"]&\ ...\ \ar[r,"\pi_1"]& A_1=\mathrm{Alb}(X)
    \end{tikzcd}
    \end{equation}

We have the following properties:
\begin{itemize}
    \item $\pi_1(A_s)=\Gamma^{s}$ and $\alpha_s$ induces the natural group homomorphism $\Gamma\to \Gamma^s$,
    \item The universal covering of $A_s$ is analytically isomorphic to a complex vector space $\C^{N_s}$.
\end{itemize}

The construction goes roughly as follows. Due to a theorem of Malcev (Theorem 6, \cite{Malcev}), for $\Gamma^s$ being a finitely generated torsion-free $s$-nilpotent group, we have a simply connected real $s$-nilpotent Lie group, say $G^{s}(\R)$, with $\Gamma^s$ as its uniform lattice. Moreover, this $G^{s}(\R)$ is unique up to isomorphism. Now take $G^s(\C)$ and we need to modula the ``antiholomorphic part'' $F^0(G^s(\C))$ (see approach 2, section 4 in \cite{Leroy}) to make sure (i) $\Gamma^{s}$ is a free uniform lattice of $G^s(\C)\slash F^0(G^s(\C))$ and (ii) $\alpha_s: X\to A_s:=\Gamma^s\backslash G^s(\C)\slash F^0(G^s(\C))$ is holomorphic. 

\subsection{Canonical coordinates for nilpotent groups}\label{subsec:canonical_coord}
This section is taken from the work of Malcev (\cite{Malcev}). For $D$ as a finitely generated torsion-free nilpotent group, we say the basis $\{d_1,\cdots, d_r\}$ of $D$ is canonical if we have 
\begin{itemize}\label{discrete_coordinate}
    \item any $d\in D$ can be written uniquely as $d=d_1^{n_1}\cdot\ldots \cdot d_r^{n_r}$ where the powers are all integers,
    \item the collection of elements in the form $d_i^{n_i}\cdot\ldots\cdot d_r^{n_r}$ form a normal subgroup $D_i\subset D$,
    \item $D_i\slash D_{i+1}$ is infinite cyclic.
\end{itemize}
and we call $\{n_1,\ldots,n_r\}$ the canonical coordinates. 

Moreover, if we embed $D$ into the simply connected real Lie group $G$ as a uniform lattice, we have a system of $1$-parameter subgroups $x_i(t)$ (with $t\in \R$ is the parameter), $1\leqslant i\leqslant r$ such that 
\begin{itemize}\label{continous_coord}
    \item for any $g\in G$, it can be written as $g=x_1(t_1)\cdot\ldots\cdot x_r(t_r)$ with $t_i\in \R$,
    \item the collection of elements in the form $x_i(t_i)\cdot\ldots\cdot x_r(t_r)$ form a normal subgroup $G_i\subset G$,
    \item $G_i\slash G_{i+1}$ is $\R$,
    \item $x_i(1)=d_i$ for $1\leqslant i \leqslant r$.
\end{itemize}

These $(t_1,\ldots,t_r)$ provide a real analytic coordinate of $G$, and they are called the coordinates of the second kind. We have the following property of the coordinates of the second kind.
\begin{proposition}(Lemma 2, \cite{Malcev})\label{prop:bounded_second_coord}
Any discrete subgroup $H\subseteq G$ is a uniform lattice if $H$ contains $x_i(1)$ for all $1\leqslant i\leqslant r$. In particular, the subset with bounded values in coordinates of the second kind is relatively compact.    
\end{proposition}

Compared with the descending central series for a $s$-nilpotent group $\Gamma$ introduced in the last section, we can write $\Gamma_i\slash \Gamma_{i+1}$ as the abelian group generated by least elements $e_{i,1},\ldots, e_{i,m_i}$, and we can arrange them together, say $\{e_{1,1},\ldots,e_{1,m_1};\ldots;e_{s,1},\ldots,e_{s,m_s}\}$ and relabel them (with order preserved) as $\{d_1,\ldots,d_r\}$. Roughly speaking, we write any element using the product from the least commutative part to the most commutative part, from left to right.

\subsection{On the coverings of a nilmanifold}\label{subsec:cover_nilmanifold}
 
Now let $M$ be the compact manifold ${G\slash D}$. This $M$ is a so-called nilmanifold, i.e. a compact manifold with a transitive action by a connected nilpotent Lie group. Now take $\hat{M}$ as any regular covering of $M$ with fundamental group $\Theta$, we have $G$ acting transitively on $\hat{M}=G\slash \Theta$ by multiplication on the left. By Theorem 2 of \cite{Malcev}, this $\hat{M}$ is a topological product $M_1\times \R^{k}$ for some integer $k$ and a nilmanifold $M_1$. We can tell that $\hat{M}$ has two ends if $k=1$ and one end if $k\geqslant 2$. We can also be very specific about this number $k$. We call a canonical generator $d_i$ with $d_i^{l}\in\Theta$ for some $l\in\Z$ a generator of type I; otherwise, we call it a generator of type II. Then $k$ is exactly the number of type II generators.

\subsection{Proof of the main theroem}
Now we are ready to prove the main theorem.

\begin{proof}[Proof of Theorem \ref{thm:nilpotent_covering}]    
Asssme that $D$ is a torsion-free $s$-nilpotent group with canonical basis $\{d_1,\ldots,d_r\}$. The given group homomorphism $\rho:\Gamma\to D$ factors through $\Gamma^s$, and we still denote the homomorphism from $\Gamma^s\to D$ as $\rho$. We are led to consider the following lift of $\alpha_s$ to the proper holomorphic map $f$ between two $\rho$-coverings:
\begin{center}
\begin{tikzcd}
{X}_{\rho} \ar[d]\ar[r,"{f}"]& M_{\rho}\ar[d]\\
X\ar[r,"\alpha_s"]& A_s=:M
\end{tikzcd}    
\end{center}

Denote the subset of type II generators as $\{d_j\}_{j\in J}$. From the argument of \cref{subsec:cover_nilmanifold}, we know the number of set $|J|\geqslant 2$. For a fixed $d_j, j\in J$, consider the natural map $M_{\rho}={G\slash\Theta}\to (G_{j}\slash G_{j+1})\cong\R$ (notation as in \cref{subsec:canonical_coord} and \cref{subsec:cover_nilmanifold}), given by the parameter $t_j$ of $1$-parameter subgroup $x_j(t_j)$. This function is known to be real analytic, and it can be expressed as the logarithmic function which is known to be a polynomial in the nilpotent case (for the details, see
section 2 of \cite{Malcev}).

Denote the composition of $f$ and $t_j$ as $f_{j}$. Recall that in the construction of the higher Albanese map in \cref{subsec:higher_alb}, we need to first complexify the real Lie group and then take its holomorphic part, thus the $f_j$ above belongs to the real or imaginary part of a holomorphic function (which is a polynomial composed with a holomorphic function) and is thus pluriharmonic. 

Set $g=\sum\limits_{j\in J}f_j^2$, which is real analytic and plurisubharmonic. Now we show $g$ is exhaustive. For a constant $c>0$ and $g\in M_{\rho}$ with $\sum\limits_{j\in J}t_j(g)^2<c$, we know that the value of $t_j$ with $j\in J$ is bounded. Besides, the value of $t_j$ with $j\in K$ for $K$ the set of the index of type I, is also bounded from the definition of $M_{\rho}=G\slash\Theta$. We know that the set of elements in $M_{\rho}$ with bounded values in $t_j$ for any $1\leqslant j\leqslant r$ is compact by \cref{prop:bounded_second_coord}, so is its pre-image by the proper map $f$ in $X_{\rho}$, i.e. $\{x\in X_{\rho}: g(x)<c\}$ is relatively compact for every $c>0$.

The rest is the same as the argument in \cref{prop:abelian_cover}. We need to analyze the degeneracy locus of the Levi form of $g$, which ends the proof of the theorem.
\end{proof}

\section{On the number of ends of Malcev coverings}
In this section, we show that the Malcve covering of any compact K\"ahler manifold $X$ has at most one end. It is well-known that any K\"ahler group can have at most one end(\cite{Gromov},\cite{ABR}), and this can be seen as a ``nilpotent'' version of that one.

We first recall the torsion-free nilpotent completion of a given finitely generated group $\Gamma$ (see Appendix A of \cite{Campana}). With the notation as in \cref{subsec:higher_alb}, we denote $\Gamma_s'=\{g\in \Gamma\colon g^n\in \Gamma_s \text{\ for some integer\ }n\}$ and take $\Gamma_{\infty}'=\bigcap\limits_{s\geqslant 1}\Gamma_{s}'$. Denote the torsion-free nilpotent completion of $\Gamma$ as $\Gamma^{\mathrm{nilp}}_0:=\Gamma\slash \Gamma_{\infty}'$ and $\rho_0^{\mathrm{nilp}}:\Gamma\to \Gamma^{\mathrm{nilp}}_0$ the natural group homomorphism. Now if $\Gamma$ is the fundamental group of a compact K\"ahler manifold, the regular covering of $X$ associated with this $\rho_0^{\mathrm{nilp}}$, denoted as $X^{\mathrm{nilp}}$, is called the Malcev covering of $X$.
The argument below is inspired by the proof of theorem 0.1 in \cite{Claudon}. We also need the following stability of the image of the series of higher Albanese maps.

\begin{lemma}[Proposition 1 in \cite{Leroy}, and Lemma 3.1 in \cite{Claudon}]\label{lem:stable_image}
With the diagram \cref{diag:higher_alb}, there exists $k\geqslant 1$ such that the normalization of the image of $X$ via $\alpha_s$ is stable when $s\geqslant k$. That is to say that if we denote $Y_s=\mathrm{Image}(\alpha_s)$ and its normalization as $Z_s$, then these $Z_s$ are biholomorphic to each other if $s\geqslant k$. 
\end{lemma}

\begin{theorem}\label{thm:one_end_malcev}
Let $X$ be any compact K\"ahler manifold with fundamental group $\Gamma$, then its Malcev covering has at most one end.
\end{theorem}

\begin{proof}
We only need to focus on the case that $\Gamma^{\mathrm{nilp}}_0$ is not finite and show that in this case $X^{\mathrm{nilp}}$ has one end. Considering the series of higher Albanese maps, by \cref{lem:stable_image}, we know that the image will be stable after finitely many steps. Denote this common image as $Y$ and its normalization as $Z$. We consider the following two cases.

\textit{Case 1}: the image of $\alpha_1$ has dimension $1$. The image $\alpha_1(X)$ is then smooth and $\alpha_1$ has connected fibers (Proposition 9.19, \cite{Ueno}). Then the smooth model $Z=\alpha_1(X)$, and by Theroem 2.2 of \cite{Campana} \footnote{This is also a result of 1-formality of K\"ahler groups\cite{DGMS}.}, we have $\Gamma_0^{\mathrm{nilp}}\cong (\pi_1(Z))_0^{\mathrm{nilp}}$. Since $Z$ is a Riemann surface, a theorem of Baumslag (\cite{Baumslag}) tells that $\pi_1(Z)_0^{\mathrm{nilp}}=\pi_1(Z)$, i.e. the Malcev covering of $Z$ is exactly the universal covering $\widetilde{Z}$ of $Z$. By the uniformization theorem, $\widetilde{Z}$ has exactly one end (for we only consider the case with infinite fundamental groups). Now consider the lift of (normalization of) the albanese map

\begin{center}
\begin{tikzcd}
X_{0}^{\mathrm{nilp}} \ar[d,swap] \ar[r,"\alpha^{\mathrm{nilp}}"]& \widetilde{Z}=Z_{0}^{\mathrm{nilp}}\ar[d]\\
X\ar[r,"\alpha"] & Z
\end{tikzcd}    
\end{center}
which is a proper surjection onto $\widetilde{Z}$ with connected fibers, thus $X_0^{\mathrm{nilp}}$ has exactly one end.

\textit{Case 2:} if the image of $\alpha_1$ is bigger than $1$, then the dimension of stable image $Y$ is $\geqslant 2$. Denote the normalization of $Y$ as $Z$ and use the same notation $\alpha_s:X\to Z$ for the normalization of $\alpha_s:X\to Y_s\cong Y$. Now we have

\begin{center}
\begin{tikzcd}
X^{\mathrm{nilp}} \ar[d,swap] \ar[r,"\alpha^{\mathrm{nilp}}"]& \hat{Z}\ar[d]\ar[r]&\hat{Y}\ar[d]\ar[r,hookrightarrow]& \widetilde{A_s(X)} \ar[d] \\
X\ar[r,"\alpha_s"] & Z \ar[r,"\beta"]& Y \ar[r,hookrightarrow] & A_s(X)
\end{tikzcd}    
\end{center}
where
\begin{itemize}
    \item the map $\beta: Z\to Y$ is a (finite) normalization of $Y$ and thus does not change the fundamental group,
    \item $\hat{Z}=Z\times_{Y}\hat{Y}$ and $X^{\mathrm{nilp}}_0=X\times_{Z}\hat{Z}$,
    \item the two coverings $\hat{Y},\hat{Z}$ of $Y, Z$ respectively are associated with the surjection $\rho_s:\pi_1(Y)\cong \pi_1(Z)\to \Gamma^s=(\Gamma\slash\Gamma_{s+1})\slash\mathrm{Tor}$ (notation as in \cref{subsec:higher_alb}),
    \item $\widetilde{A_s(X)}$ is the universal covering of $A_s(X)$, which is a complex vector space.
\end{itemize} 
Due to the subjectivity of $\rho_s$, the $\hat{Y}$ embedded in a complex vector space $\widetilde{A_s(X)}$ is closed and connected, thus Stein. Then $\hat{Z}$, which admits a finite ramified covering over $\hat{Y}$, is also Stein. Notice that the proper morphism $\alpha^{\mathrm{nilp}}$ might have unconnected fibers. In this case, we take its Stein factorization, say $X^{\mathrm{nilp}}\xrightarrow{s}S\xrightarrow{t}\hat{Z}$ such that $t$ is finite and $s$ has connected fibers. Then $S$, as a finite covering over a Stein manifold $\hat{Z}$, is itself connected and Stein. Since $\mathrm{dim}(S)\geqslant 2$, we know that $S$ has one end (see Corollary 4.10 of \cite{BuanicuaStuanuac}). Then $X^{\mathrm{nilp}}$ has one end since $s: X^{\mathrm{nilp}}\to S$ is a proper surjection with connected fibers.
\end{proof}

\begin{remark}\label{rmk:compare_known_results}
The combination of \cref{thm:one_end_malcev} and \cref{thm:nilpotent_covering} tells us that the Malcev covering of any compact K\"ahler surface is holomorphically convex. This is certainly not new and holds in any dimension (i.e. not necessary for surfaces). The proof was given by Katzarkov in the projective case (\cite{Katzarkov97}) and by Leroy in the K\"ahler case (\cite{LeroyThesis} and \cite{Claudon}). Thus the main theroem is new only for intermediate coverings over compact K\"ahler surfaces.  
\end{remark}

\section*{Acknowledgments}
I would like to thank Professor Mohan Ramachandran for communicating the proof of \cref{prop:abelian_cover} to me. The author is partially supported by the China Postdoctoral Science Foundation 2023M743396.

\bibliographystyle{ssmfalpha} 
\bibliography{main}

\end{document}